### v1

# Semantic Trajectory Data Mining with LLM-Informed POI Classification

Abstract
Human travel trajectory mining is crucial for transportation systems, enhancing route optimization, traffic management, and the study of human travel patterns. Previous rule-based approaches without the integration of semantic information show a limitation in both efficiency and accuracy. Semantic information, such as activity types inferred from Points of Interest (POI) data, can significantly enhance the quality of trajectory mining. However, integrating these insights is challenging, as many POIs have incomplete feature information, and current learning-based POI algorithms require the integrity of datasets to do the classification. In this paper, we introduce a novel pipeline for human travel trajectory mining. Our approach first leverages the strong inferential and comprehension capabilities of large language models (LLMs) to annotate POI with activity types and then uses a Bayesian-based algorithm to infer activity for each stay point in a trajectory. In our evaluation using the OpenStreetMap (OSM) POI dataset, our approach achieves a 93.4% accuracy and a 96.1% F-1 score in POI classification, and a 91.7% accuracy with a 92.3% F-1 score in activity inference.
I Introduction
The rapid development of Internet-of-Thing (IoT) technology and applications facilitates the evolution of intelligent transportation systems and brings a new era of data collection [1, 2, 3, 4, 5]. This flourish yields an abundance of trajectory data from a wide range of connected devices and allows us to explore human travel behavior with additional detail and accuracy. To be specific, the advent of GPS-enabled mobile devices has revolutionized the tracking of individuals, objects, vehicles, and even animals through the collection of digital traces or trajectories [6, 7, 8, 9]. Furthermore, integrating these location trajectories with semantic information from spatial data sources like Points of Interest (POI) datasets enriches our understanding of the behavior of moving entities. This integration facilitates applications in traffic management, epidemic tracking, and human movement studies [10, 11, 12].
To deepen the analysis of human travel trajectories for studying behavior patterns, it is important to comprehend the meaning behind each movement or stay point, which is aggregated based on GPS coordinates. This depth of analysis is achieved by incorporating POI data, where each stay point is linked to nearby POIs to infer the potential activities (e.g., work, exercise, and religious activity) at each location. Activity-based models (ABM) are established by providing insights into why people travel to specific locations. This data mining method allows us to add a layer of intention and purpose to the trajectories, deepening our comprehension of the underlying human behaviors that drive travel patterns. Therefore, the trajectory mining approach, which involves classifying POIs with activity types (interchangeably referred to as activity categories) and inferring activities at each stay point, is pivotal in translating raw data into meaningful insights that can inform various societal and technological applications.
The effectiveness of the trajectory mining process significantly depends on the accuracy of the POI classification. POI data play a vital role in numerous applications including navigation solutions, local search functionalities, smart road services, and human travel pattern studies [13, 14, 15, 16, 17]. A reliable POI classification algorithm can notably enhance the usability, functionality, and effectiveness of these applications by ensuring the quality of POI data categorization results. Despite their importance, existing learning-based classification methods suffer from considerable variability within open-source datasets like OpenStreetMap (OSM) [18], and discrepancies widen when multiple data sources are considered. This inconsistency not only complicates schema alignment but also significantly limits the amount and quality of usable data. Efforts to synchronize these disparate data schema are time-consuming and often result in loss of data granularity and reliability.
To address these challenges, we introduce a novel framework that leverages the strong inferential and comprehension capabilities of Large Language Models (LLMs) for POI classification. Once POIs are accurately classified with activity types, the next step is to infer the likely activity type occurring at each stay point using a Bayesian-based algorithm. The proposed approach is validated on the OSM dataset, demonstrating superior performance.
Compared to existing literature, the main contributions of this paper are the following:
  • We introduce a novel trajectory data mining framework that integrates LLM-based POI classification with the Bayesian-based activity inference algorithm. The outcome of this data mining effort, annotated trajectories, enhances the understanding of human travel patterns while ensuring the privacy of individuals during research. To the best of our knowledge, this is the first attempt to apply large language models to solve the POI classification problem.
  • Our proposed framework is highly adaptable and can be easily deployed to POI or trajectory datasets in different regions without the need for additional training. The POI classification method does not require high data integrity and is suitable for open-source datasets with incomplete data.
  • Leveraging the semantic information inferred from POIs provided by LLM outputs, our activity inference algorithm achieves precise point-level inference.
II Literature Review
II-A POI Classification
The classification of POIs involves systematically matching each POI observation into defined categories. These categories can be based on the type of location, for instance, restaurants, schools, or bars, or the type of activity associated with the location, such as shopping, working, or exercising.
Learning-based methods like support vector machines (SVM) [19] were first applied for the POI classification task. By incorporating additional textual information surrounding a POI’s name in its context, 73% accuracy was achieved on the Yelp dataset with 20 different categories [20]. Besides, combining the information from neighborhood POIs based on coordinate with additional features extracted from POIs’ name and description [21] reached an average 70% accuracy and 60% F-1 score in multiple cities for the Yelp Datasets with 10 categories.
Spatial, textual, and property features were also integrated to increase the performance of POI classification [22]. In the OSM dataset with 14 categories, the SVM-based method achieved 60% accuracy, while a clustering approach coupled with k-Nearest Neighbors (kNN) attained a 91% top-10 accuracy. Additionally, search engines, which have access to real-time data, have been employed to extend original POI datasets [23]. In the POI dataset with 10 categories from Beijing, 86% accuracy and 40% F-1 score were reached by using attention-based convolutional neural networks.
As the number of extended features grows, the investigation of feature importance in classification outcomes has become pertinent. A variety of popular learning-based methods, including Linear Discriminant Analysis, Support Vector Machines (SVM), K-Nearest Neighbor (kNN), Random Forest (RF), and eXtreme Gradient Boosting (XGB), have been applied to determine the significance of different features in classification tasks [24]. Research has demonstrated that the importance of a feature can vary according to geographical location and the categories being predicted, which complicates the development of a universally high-performing algorithm for diverse datasets.
Previous approaches mainly focused on learning-based algorithms and attempted to increase classification performance by feature extending or additional information from external data sources. However, in real-world scenarios, missing features are common, and it can be either difficult or time-consuming to gather and synchronize the necessary data for training from external sources. Maintaining data consistency across different sources adds another layer of complexity. Furthermore, arbitrary labeling styles can complicate classification efforts. For instance, the OSM data points from Egypt frequently feature names solely in Arabic, without English translations, limiting the effectiveness of classification algorithms that are not designed to handle multilingual contexts.
The bottlenecks mentioned above highlight the necessity for a more robust and adaptable algorithm capable of effectively processing and integrating diverse and incomplete POI datasets. Recently, large language models trained by trillion-level tokens and diverse data sources including internet content have shown their superiority in various tasks like chatbots, robot control, law counseling, and even code writing [25, 26, 27, 28]. Benefiting from the transformer architecture, and diverse training data, LLMs also show a strong generalization and inferring ability when facing unseen scenarios [29]. In our work, we address these existing challenges by introducing an LLM-based approach. This method enables effective classification across different datasets without the necessity for integrity data or additional feature extension.
II-B Activity Inference
Figure 1: Architecture of the proposed framework from POI classification (left) to activity Inference (right).
Trajectory data that is collected using GPS-integrated devices typically consists of sequences of geo-coordinates documenting daily trips. Although these data primarily record movement, they implicitly reveal significant insights into human activities, such as purchasing goods or dining out. However, inferring the purpose of trips is challenging due to the inaccuracy of GPS data in pinpointing the visited POIs and the incomplete information from POI datasets in explaining activity purposes [30].
Previous studies have extracted spatial and temporal features from movement trajectories to represent characteristics of travel behaviors for activity inference [31]. Spatial patterns include metrics like visit frequencies and radius of gyration, while temporal patterns focus on the duration and type of day. Rule-based methods, which utilize these features, have shown high accuracy in inferring mandatory activities such as home and work, due to their strong periodicity and higher visit frequency [32].
However, these methods fall short in inferring more complex non-mandatory activities like eating out or visiting friends. To enhance context in activity inference, different methods have integrated POI datasets. Graph Neural Network (GNN) methods, for instance, construct a weighted directed graph by considering the POI distribution within an activity zone and connecting regular areas to build the graph [33]. The functionality and popularity of POIs in trip end areas have also been considered in Dynamic Bayesian Network models. The model’s performance is evaluated in predicting the correct activity type as the top choice or within the top 3 choices, and DBN could achieve 64.1% at rank 1 and 87.8% at rank 3 [34]. These models have improved activity inference for non-mandatory types, but the information from POIs remains imprecise and aggregated zonal information may diminish the utility of precise GPS trajectory data.
Additionally, researchers have examined the sequential properties of activities, noting the importance of temporal periodicity [34]. Prior knowledge of these patterns, obtainable from datasets like household travel surveys, is crucial yet underutilized in current research, which often overlooks the temporal distribution of activities.
To address these issues, our model is designed to accurately infer activity types using point-level POIs, rather than relying on broader zonal-level data, allowing for more precise geographic inference. Furthermore, prior knowledge of activity temporal distributions significantly enhances inference accuracy. We address these challenges through the introduction of a Bayesian-based algorithm that considers the nearby POI context.
III Methodology
The architecture of our trajectory mining framework is illustrated in Fig. 1. We first perform POI classification by reformulating the task description and candidate activity types into natural language representations. This information is then combined with the POI observations to determine the three most likely activity types for each POI using LLM. Meanwhile, we engage in prompt engineering to refine the quality of outputs based on the model’s performance. Following this, we integrate stay point  [35] information and implement a Bayesian-based inference algorithm to accurately associate each stay point with potential POIs and activity types in the human trajectory data.
III-A Problem Formulation
Given the POIs and stay points from trajectory data, the task is to infer activity types for stay points in the trajectory data, and the mathematical expressions and definitions are provided:
POI: The point of interest observations are denoted as $\mathcal{P}=\{P_{1},P_{2},...,P_{I}\}$, where each observation $P_{i}$ has $\langle\text{Name},\text{Lon},\text{Lat},f_{1},f_{2}...f_{n}\rangle$, indicating the name, longitude, latitude and available features($f_{n}$) of the POI
Trajectory: The trajectory data traces are denoted as $\mathcal{T}=\{Traj_{1},Traj_{2},...,Traj_{N}\}$, where $Traj_{j}$ denotes the trajectory for the $j$-th person.
Stay Point: For each trajectory of person $n$, the stay points are defined as any location where the person remains stationary for a duration of time. Thus, a trajectory could have multiple stay points $Traj_{j}=\{S_{n1},S_{n2},...,S_{nK}\}$, where $S_{nk}$ denotes the $k$-th stay point of person $n$. Each stay point has information of $\langle T_{S},T_{E},\text{Lon},\text{Lat}\rangle$, representing the start time, end time, longitude, and latitude of the stay point.
Given trajectory $Traj_{n}$, information of the stay points $S_{nk}$ $\langle T_{S},T_{E},\text{Lon},\text{Lat}\rangle$, and POI records $\mathcal{P}$, the goal is to predict top three activity denoted as $T$(POI Classification), and the corresponding probabilities denoted as $P$ from $\mathcal{P}$ and inference the activity type $T_{nk}$ for each stay point(Activity Inference).
III-B POI Classification
Figure 2: Example of the prompt and output from LLM.
Different from learning-based algorithms that derive relationships between POI features and categories from labeled data samples, LLM leverages its powerful inferencing capabilities to respond to open-ended text-based questions without requiring additional training. In our approach, we reframe the classification task into a text-based question format where LLMs excel and design efficient prompts that precisely describe the problem, allowing the LLM to interpret and classify POIs effectively based on the context provided in the prompts.
As shown in Fig. 3 as an example, the prompt is composed of three distinct components:
Task Description: This component outlines the POI classification task, specifying the input data, the types of information included, and the desired output. It sets the framework for how the LLM should approach data analysis and classification.
Category Description: This component provides a comprehensive description of each candidate category we are interested, including definitions and examples if necessary. This detailed categorization helps guide the LLM in accurately classifying the POIs into relevant categories.
POI Observation: In this part, the features from the dataset and their corresponding values are represented in natural language format, ensuring the model will comprehensively understand and process the characteristics of each POI as described in the data.
Notably, prompt fine-tuning can be utilized to meet specific needs when working with diverse datasets or various category sets, as LLMs are not limited by text input formats. For instance, in the Egyptian POI collection where some POI names are provided only in Arabic, we can enhance the model’s understanding by explicitly stating in the prompt like “some names in the ‘name’ column are in Arabic.” Additionally, for datasets that only contain POIs from public spaces, we can point out the source of the dataset in the prompt, and amenities like “toilets” will not be classified as “home activity.” Moreover, when dealing with real-world datasets where features often missing such as a POI labeled only with the amenity type “visitor parking”, it can be challenging to determine whether this POI belongs to the category “visit friends or relatives” or “pick up/drop off somebody.” For such cases, we can refine the prompt to guide the LLM in identifying the three most relevant categories for the POI, complete with probability scores for each. This tailored approach ensures that the model provides meaningful classifications even with limited data, and these probabilities can be further utilized in other tasks, such as activity inference.
III-C Activity Inference
Activity inference involves considering many factors, such as the time of the day, the category of the visited location, and the functionality of the location, requiring detailed information. However, the privacy implications of the use of location-based services [36], especially identifying specific POIs in the trajectory, are raising serious concerns. It could potentially disclose any specific user’s daily activities, interests, and even personal habits. Thus, it is not appropriate to match the stay points in real trajectory data with specific POIs to conduct the inference. Therefore, an effective activity inference model must carefully balance the inclusion of detailed information with privacy safeguards. In the following parts, we elaborate a rule-based algorithm for mandatory activities inference and a probabilistic model for non-mandatory activities inference as shown in Fig. 3.
Figure 3: Flow chart of activity inference procedure.
Mandatory Activity Inference: Our mandatory activity inference algorithm is designed to identify three primary activities: Home, Work, and School, inspired by a rule-based approach [32]. Firstly, we determine the Home activity by selecting the stay point with the highest visit frequency during both weekends and weekdays between 7 pm and 8 am. Next, to infer the Work activity, we identify stay points not previously labeled as Home and calculate the total distance traveled from Home to each stay point during weekdays between 8 am and 7 pm. The work location is determined by identifying the stay point with the highest frequency of visits during work hours and the greatest total distance traveled from home. Finally, for inferring the School activity, we identify stay points with the highest total distance traveled from Home during weekdays between 8 am and 7 pm. If a stay point is within 20 meters of an ‘education’ POI, it is labeled as the School location. This rule-based algorithm effectively assigns mandatory activities based on stay point characteristics and geographical proximity to relevant POIs.
Non-mandatory Activity Inference: Non-mandatory activities demonstrate different patterns from mandatory activities in the visit periodicity and frequency. Thus, we apply a probabilistic framework to predict activity types $T$ based on location $L$ and start time $S$ using a combination of the law of total probability and Bayes’ theorem. We begin by defining the probability of an activity type $T$ given a location $L$ and start time $S$ using the law of total probability:
$P(T\mid L,S)=\sum_{i}P(T\mid L_{i},S)P(L_{i})$
Next, we refine $P(T\mid L_{i},S)$ using Bayes’ theorem:
$P(T\mid L_{i},S)=\frac{P(S\mid T)P(T\mid L_{i})}{P(S\mid L_{i})}$
Where $P(S\mid L_{i})$ is simplified under the assumption of independence between $S$ and $L_{i}$:
$P(S\mid L_{i})=\sum_{T}P(S\mid T)P(T\mid L_{i})$
This leads us to the final expression that calculates the probability of a certain activity type:
$P(T\mid L,S)=\sum_{i}\frac{P(S\mid T)P(T\mid L_{i})}{P(S\mid L_{i})}P(L_{i})$
This methodology allows for an integrated analysis of the effects of location and time on activity types, providing a robust framework suitable for use in location data analytics to predict varying activities effectively.
IV Experiment
IV-A Dataset
To evaluate our approach across various contexts and assess its effectiveness, we use open-source data from Los Angeles (LA) County and Egypt, sourced from OSM. This data enables us to rigorously test the performance of our POI classification algorithm.
The dataset is structured as a directed, connected network with POIs, each associated with metadata tags found on the network’s edges or vertices. Given that OSM is a free, open geographic database maintained by volunteers through open collaboration, the data exhibits significant diversity in terms of quality, format, and variety. This includes a range of location-specific details such as buildings, amenities, and infrastructure elements like roads and railways. While it offers higher information entropy to the LLM, it also complicates predictions due to ambiguous or contradictory labels. For instance, a church in OSM could be labeled as a place of worship, a tourist attraction, and an office building simultaneously. While this might seem contradictory at first glance, such a POI is applicable in all three contexts, whether an agent is going for worship, tourism, or work.
Despite the crucial “names” feature being present, 89% of the content in other features such as “amenity,” “building type,” and “land use” is missing in the Egypt dataset, and similarly, 91% is missing in the LA dataset. Utilizing datasets like OSM enables us to evaluate the resilience of our methodology in challenging conditions due to the inconsistent labeling style and incomplete data in the real-world scenario.
IV-B Evaluation Setting
Our study covers 15 activity categories [37] to capture a wide range of daily activities, ensuring the dataset reflects the diversity of human behavior, as detailed in Table I.
IV-B1 POI classification
Evaluating the performance of our POI classification method presents challenges typical of an unsupervised learning setting, where no pre-existing ground truth is available. To address this, we randomly sample 500 observations from each POI dataset in both Egypt and LA County and create a manual ground truth. The accuracy of our method was then benchmarked against this human-labeled ground truth. To minimize labeling bias, five volunteers are enlisted to categorize the labels under uniform standards. Finally, we employ the language models ‘gpt-3.5-turbo (gpt3.5)’ and ‘gpt-4 (gpt4)’ from OpenAI, prompting them to identify the top three most relevant categories for each POI, along with the associated probabilities.
TABLE I: Activity category code and their corresponding descriptions
Category Description
1 Home activities (sleep, chores, etc) or Work from home
2 Work-related activity or Volunteer
3 Attend school
4 Attend child or adult care
5 Buy goods (groceries, clothes, gas)
6 Buy services (dry cleaners, banking, service a car)
7 Buy meals (go out for a meal, food, carry-out)
8 General errands (post office, library)
9 Recreational activities (visit parks, movies, bars)
10 Exercise (jog/walk, walk the dog, gym, etc)
11 Visit friends or relatives
12 Health care visit (medical, dental, therapy)
13 Religious or community activities
14 Something else
15 Drop off/pick up someone
IV-B2 Activity Inference
After completing the POI classification using the ‘gpt-3.5’ model, we obtain a dataset of 85,696 POIs, each annotated with possible activity categories, from LA County. This dataset serves as the input for the activity inference process, as illustrated in Fig. 1.
To assess the efficacy of our proposed activity inference model, we curate a test trajectory dataset by extracting activities from the 2017 National Household Travel Survey (NHTS) California Add-on dataset [37], focusing on LA County. This dataset provides detailed daily travel trajectories, with each trajectory consisting of activities and a rough location at the zonal level. However, due to the lack of POIs associated with these activities, we utilize LLM for preliminary POI selection from the rough location zone, followed by manual curation to rectify any erroneous assignments. It is important to note the limitations of GPS-enabled smartphones, which typically provide accurate location data within a 4.9-meter radius under ideal conditions [38]. However, accuracy diminishes in urban environments near structures like buildings, bridges, and dense foliage. To evaluate the model’s performance and robustness against real-world conditions, we introduce noise at three levels: 5 meters, 10 meters, and 20 meters. The noise, simulated using a Gaussian kernel, is added to the latitude and longitude coordinates of the POIs. Subsequent to this augmentation, our test dataset comprises 362 individuals and 2,007 activities, consisting of 1,724 mandatory activities and 283 non-mandatory activities.
IV-C Result
IV-C1 POI Classification
Figure 4: Example of POIs and classification results from (a) LA county and (b) Egypt.
Considering there is a significant amount of incomplete features in the dataset. For observations that only have limited features available, such as a POI with only a name labeled like “Volkswagen”, it’s hard to indicate whether the individual is there for car purchasing (category 5), auto maintenance (category 6), or is simply an employee at the location (category 2). As a result, in the calculation of overall accuracy and macro F-1 score, we consider the classification correct if one of the three predicted activities from the model matches the manually labeled activity category. Additionally, we calculate the hit rate for each position of the predicted activity code, recorded as “Hit@n”. The model returns these codes in decreasing order of probability, as specified in the prompt.
TABLE II: Performance metrics for POI classification with different models.
Metric Egypt(gpt3.5) Egypt(gpt4) LA(gpt3.5) LA(gpt4)
Accuracy 90.3% 93.4% 82.5% 91.4%
Hit @1 61.6% 74.7% 65.2% 75.7%
Hit @2 22.6% 9.7% 15.4% 8.7%
Hit @3 6.1% 9.0% 1.9% 7.0%
F-1 Score 91.50% 96.10% 82.1% 92.9%
As shown in Table II, our model achieves a 93.4% accuracy for the Egypt dataset and 91.4% accuracy for the LA county dataset, with F-1 scores of 96.10% and 92.9% respectively when using the gpt4 model. This demonstrates its superiority and robustness compared with previous learning-based methods.
Compared to gpt3.5, the more advanced gpt4 model outperforms in both accuracy and F-1 score, especially for challenging datasets like LA country where more feature contents are missing. Besides, the analysis of the hit rate of the three predicted categories shows that while the “Hit @1” rate dominates when using both models, the “Hit @1” rate of gpt4 is approximately 10-13% higher than that of gpt3.5. This improvement underscores the advanced model’s enhanced comprehension abilities, leading to more robust performance in classification.
Some instances of the classification results are listed in Fig. 4. For POI 1, with only the name “KFC”, the model confidently categorized category 7 (Buy meals) as its first match with a probability of 0.7, and the semantically related “buy” category 5 (Buy goods) as the second choice. For POI 2 with a clear description, the model predicted category 10 (Exercise) as the best match with a 0.7 probability, and category 9 (Recreational activities) as the next likely category with a 0.2 probability. In the case of POI 3, a cafe, the model assigned close probabilities between categories 7 (Buy meals) and 9 (Recreational), since a cafe usually includes both the dining and recreational social attributes. For POI 4 and POI 5 whose names were documented in Arabic(“Roman theater” and “Dr. Rafiq Suleiman’s clinic”), the model successfully understood the meaning and labeled them with corresponding categories. Besides, 14 (Something else) was chosen as the complementary third choice in all cases above as there are no other related categories. These impressing instances again demonstrate the strong comprehension ability of our approach, offering logically predicted probabilities that enhance subsequent activity inference tasks.
IV-C2 Activity Inference
In evaluating our activity inference model on the test dataset, we utilize accuracy @1, @2, and @3 under 5m, 10m, and 20m levels of noise as evaluation metrics. These metrics measure the model’s performance in predicting activity types for stay points. Accuracy @1 represents the proportion of stay points where the correct activity type is the top prediction, while accuracy @2 and @3 indicate the percentages where the correct activity type falls within the top two and top three predictions, respectively.
Figure 5: Semantic trajectories examples after data mining. (a) Activity-annotated stay points of trajectories in LA. (b) Activity timelines of trajectories.
According to Table III, which outlines the performance metrics for activity inference under varying levels of noise, the algorithm maintains a commendable degree of stability even as the standard deviation (SD) of noise increases. Specifically, for non-mandatory activities, Acc @3 remains above 80%, indicating a resilient performance against the perturbations caused by noise. With a 5-meter standard deviation in noise, the accuracy at Acc @3 is recorded at 88.4%, which diminishes to 84.2% when the noise level is increased to 20 meters. This reduction, while notable, is not drastic, supporting the assertion that the model upholds a stable predictive capability in the face of increasing noise levels. Particularly with respect to GPS trajectory datasets, the model demonstrates robust inference capabilities at a 5-meter noise accuracy. These findings not only validate the stability of the model but also highlight its potential for effective deployment in real-world scenarios where varying degrees of GPS accuracy are a common occurrence.
TABLE III: Performance metrics for non-mandatory activity inference under different levels of noise
SD of noise Type Acc @1 Type Acc @2 Type Acc @3
5m 75.0% 84.1% 88.4%
10m 73.6% 81.8% 85.3%
20m 71.9% 80.5% 84.2%
As shown in Table IV, the results indicate that mandatory activities (categories 1-3) exhibit consistent performance due to the deterministic nature of their patterns. On the other hand, activities characterized by more complex and variable patterns present greater challenges for prediction. Despite this, categories 4, 5, 7, 8, 12, and 13 demonstrate commendable accuracy. Notably, categories such as 5 (purchasing goods), 7 (dining out), 12 (availing services), and 13 (religious activities) show robust performance, which can be attributed to the ample instances available within the test dataset.
Conversely, the categories of 11 (visiting friends or relatives), 14 (miscellaneous activities), and 15 (transporting others) prove to be more elusive in the context of stay point detection. These activities are less associated with consistent POIs, leading to relatively poorer model performance. This suggests that the rule-based algorithms are effective for activities with clear-cut patterns.
TABLE IV: Performance metrics for activity inference under 5m noise for different activity types
Types Acc@1 F1 score Types Acc@1 F1 score
1 98.3% 98.3% 9 63.1% 74.2%
2 87.2% 95.1% 10 64.7% 76.1%
3 90.0% 95.0% 11 52.2% 67.3%
4 100% 98.3% 12 100% 97.2%
5 78.5% 76.3% 13 100% 100%
6 66.7% 76.2% 14 46.2% 67.2%
7 77.4% 87.1% 15 50.1% 70.6%
8 100% 96.6% Average 91.7% 92.3%
Finally, the outcome of the proposed framework is the semantic trajectory, as illustrated in Fig. 5, where each stay point in a trajectory is annotated with the most likely activity. For instance, as shown in Fig. 5 (a), individuals may engage in activities such as commuting between their home and work locations, exercising in the morning, or purchasing goods while en route from work to home. By augmenting stay point data with activity type information, we enhance our understanding of human travel mobility and behavior patterns. Through extensive trajectory mining, we progress from raw GPS data (aggregated to stay points) to refined activity chains, marking a significant advancement in the study of human travel behaviors.
V Conclusion and Future Work
Our study has introduced a novel LLM-based framework for POI classification and activity inference, demonstrating notable improvements over traditional learning-based methods in handling incomplete and varied datasets like those from OpenStreetMap. Our approach, by leveraging the robust inferential capabilities of large language models, has shown exceptional performance with high accuracy and F-1 scores in real-world scenarios across diverse geographic locations.
Looking ahead, we aim to enhance the accuracy and robustness of POI classification methods with more sophisticated prompt engineering or the development of a fine-tuned LLM designed specifically for the POI classification task.
VI Acknowledgement
We want to thank Xiangnan Zhang and Zhaoyi Ye for their contribution to dataset preprocessing.